OpenStudy (anonymous):

The max velocity of the emitted electron is ________.,if the light of wave length 4300A is incident on the surface of potassium whoes work funtion is 2.3eV-> help please

OpenStudy (anonymous):

The Einstein equation is \[ K=h\nu - \Phi\] where K is the kinetic energy of an emitted electron, h nu is the energy of a photon (Planck's constant times the frequency of a photon), and phi is the work function. K can be rewritten as \[K=\frac{1}{2}mv^2\] so that you can solve directly for the velocity. Does that help?

OpenStudy (anonymous):

Remember that \[1 A = 1 x 10^{-10}m\] and that the frequency and wavelength of a photon are related by \[c=\nu \lambda\] \[ \nu = \text{ frequency}\] \[ c = \text{ speed of light} = 3.0x10^8m/s\] \[ \lambda= \text{ wavelength in m}\] \[ 1eV = 1.602x10^{-19}J\]

OpenStudy (anonymous):

so the velocity of the electron would be \[v=\sqrt{2\frac{h \nu - \Phi}{m_e}}\] \[h \propto \ m^2 \ kg/s \] \[\Phi \propto J=kg \ m^2/s^2\] \[m_e = \text{ mass of an electron}\]

OpenStudy (anonymous):

I guess the first thing I should have asked is if you knew what a work function was :P Sorry :(

OpenStudy (anonymous):

oooppss i guess i know ......any ways thanks for the help ... really apreciate it ....

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

@shkrina ?

OpenStudy (anonymous):

Or did I just blow up a Greek alphabet and nonsense fell out?

OpenStudy (anonymous):

sir .. thanks for ur help .. i did not get ur above .. question .. waht did not make sense for u .. well the above statement says was i know .. that procedure which ugave me .. and in my text ist diffrent .. soo just wanted 2 clarify .. ..any more questions ..sir..^_^

OpenStudy (anonymous):

@AllTehMaffs

OpenStudy (anonymous):

I just couldn't tell what your ellipses meant is all ^_^ If you understand it then I don't have any questions; I just wanted to make sure!

OpenStudy (kittiwitti1):

(Curse you typos!) @shkrina He can't see your replies or tags 'cause right now the notifications system that tells you if you've been tagged is down ^^

OpenStudy (anonymous):

thnku guys .. cyaaa ..

OpenStudy (anonymous):

ciao ^_^

OpenStudy (kittiwitti1):

Oh, don't thank me @shkrina, I did nothing ^^ Just close the question and all is done (; --- Well, and give @AllTehMaffs a medal for his help.